In a year, Seema earns ₹ 1,50,000 and saves ₹ 50,000. Find the ratio of
Money that Seema earns to the money she saves

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Money earned by Seema=₹150000
Money saved by her =₹50000
Money spent by her =₹150000 -₹50000 =₹100000
Ratio of money earned to money saved = 150000 / 50000 = 15 / 5
= 3 / 1

Thanks